• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 01:17:43 on October19,2009

    Entry number 295594

    keyword=Compton HV tripped.

    The Compton HV tripped at 01:13.

    Laser on rate was 800-900 just before the trip.

    Reset HV.

    Laser on rate after the last HV reset was ~700-750.
    Laser off rate was ~100.

    FIGURE 1